OpenStudy (anonymous):

derivative of cos^2(sin(x))

OpenStudy (anonymous):

what's the first step?

OpenStudy (anonymous):

use chain rule

OpenStudy (anonymous):

i know chain, but...

OpenStudy (anonymous):

i think im suppose to use law of logarithms to bring the square down

OpenStudy (anonymous):

why do you want to bring down the square take f(x)=sinx and g=cos^2(f) ow find g'

OpenStudy (anonymous):

-2cos(sinx) * sin(sinx)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[ f(x)= \cos^2(\sin x)\\ f'(x)=-2\cos(\sin x)\cdot\sin(\sin x)\cdot \cos x \]

OpenStudy (anonymous):

why that last cosx?

OpenStudy (anonymous):

You have to use the chain rule twice. You have basically f(x)=x^2, g(x)=cos(x), h(x)=sin(x), and you have f(g(h(x))), so apply chain rule twice.

OpenStudy (anonymous):

ty